Solve the system of equations using the addition (elimination) method.
If the answer is a unique solution, present it as an ordered pair: (x, y). If not, specify whether the answer is "no solution" or "infinitely many solutions.
x + 5y = 4
2x – 5y = 2

Answer by Alan3354(69443)

x + 5y = 4
2x – 5y = 2
-------------- Add them
3x + 0y = 6
3x = 6
x = 2
Substitute for x in either eqn
2 + 5y = 4
5y = 2
y = 2/5